AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  DESPEC  ELOG logo
Entry  Mon Oct 27 11:55:00 2025, JB, MP, GB, NH, Changing ground setup from 25102025 9x
    Reply  Mon Oct 27 18:08:11 2025, TD, Changing ground setup from 25102025 
Message ID: 727     Entry time: Mon Oct 27 11:55:00 2025     Reply to this: 728
Author: JB, MP, GB, NH 
Subject: Changing ground setup from 25102025 
12:55 We are going to try the follow changes to the AIDA scheme of 241012025.

- Fitting LK1 (aida01, aida14), removing LK1 (aida03, aida07, aida11, aida15) - only 1x LK1 needs to be fitted per Si wafer
- Removing LK4 to see what (if anything) changes.

Check that all of the screws securing the copper foil to ground are tight - snout install/removal tends to loosen screws.

- Remove the bias line between aidas 12-3-15 and aidas 16-7-11
- Remove the test circuit
 (These two should eliminate any extra ground path on the front, except via the DSSD)

- Use a multimeter to confirm 0 Ohm between all the copper foils  (short the two probes to correct for resistance of the probe leads)

- Try combinations of LK2/LK4 on the n+n fees

13:01 We check first the grounding with a multimeter and the different copper grounding components. They should all be grounded to the large iron frame, which is in turn grounded to the platform. We also checked that the CAEN HV PSU was also grounded to the platform (it is).

13:22 We removed the LK on aida03, aida07, aida11, aida15 and added LK1 to aida01 and aida14. We try starting the DAQ and seeing the response from these changes.

13:25 We tried biasing with this configuration and the first DSSD on HV01 read a leakage current of 40 uA (too high). So we biased it down and reassessed. 

13:30 It turns out we had LK1 fitted to aida04. We tried biasing again but this time DSSSD on HV01 tripped. They are fitted with LK4 so we biased down and removed these link on aida06 and aida08.

13:35 Trying to bias now with LK4 removed from aida06 and aida08. We tried fitting and removing again. Now the setup appears to bias. This was probably due to a stray piece of copper shielding touching the PCB.

14:22 We are checking the noise conditions now. TEMPS OK. FADC Align and whiterabbit timestamps OK.

14:30 With the changes above we have noticed an intense 5/2 ms pulsed signal in the wave forms. See attachment 2

15:20 We decabled everything and the current scheme from 24102025, see prev. elog entry is in attachment 1.

15:34 We have unbiased AIDA1 and the wave forms improve for AIDA2 (?), 5/2 ms pulsed signal is still observed in AIDA1
Attachment 1: AIDA_scheme_241020525(1).png  24 kB  Uploaded Mon Oct 27 14:20:25 2025  | Hide | Hide all
AIDA_scheme_241020525(1).png
Attachment 2: Screenshot_from_2025-10-27_15-28-21.png  176 kB  Uploaded Mon Oct 27 14:33:39 2025  | Hide | Hide all
Screenshot_from_2025-10-27_15-28-21.png
Attachment 3: Screenshot_from_2025-10-27_14-44-36.png  174 kB  Uploaded Mon Oct 27 14:33:39 2025  | Hide | Hide all
Screenshot_from_2025-10-27_14-44-36.png
Attachment 4: Screenshot_from_2025-10-27_14-42-57.png  162 kB  Uploaded Mon Oct 27 14:33:39 2025  | Hide | Hide all
Screenshot_from_2025-10-27_14-42-57.png
Attachment 5: Screenshot_from_2025-10-27_14-38-24.png  130 kB  Uploaded Mon Oct 27 14:33:39 2025  | Hide | Hide all
Screenshot_from_2025-10-27_14-38-24.png
Attachment 6: Screenshot_from_2025-10-27_14-37-50.png  225 kB  Uploaded Mon Oct 27 14:33:39 2025  | Hide | Hide all
Screenshot_from_2025-10-27_14-37-50.png
Attachment 7: Screenshot_from_2025-10-27_14-26-41.png  161 kB  Uploaded Mon Oct 27 14:33:39 2025  | Hide | Hide all
Screenshot_from_2025-10-27_14-26-41.png
Attachment 8: Screenshot_from_2025-10-27_14-26-28.png  898 kB  Uploaded Mon Oct 27 14:33:39 2025  | Hide | Hide all
Screenshot_from_2025-10-27_14-26-28.png
Attachment 9: Screenshot_from_2025-10-27_15-34-01.png  182 kB  Uploaded Mon Oct 27 14:34:13 2025  | Hide | Hide all
Screenshot_from_2025-10-27_15-34-01.png
ELOG V3.1.3-7933898